OpenStudy (emily_98):

What is the discriminant of the quadratic equation 3 – 4x = –6x^2?

OpenStudy (welshfella):

rewrite it as follows 6x^2 - 4x + 3 = 0 now compare it with the general form;- ax^2 + bx + c = 0 so a = 6 and what is b and c?

OpenStudy (welshfella):

the discriminant is b^2 - 4ac just plug in values of a,b and c and work it out

OpenStudy (anonymous):

\[D = b ^ 2 - 4ac \] in the @welshfella equation

OpenStudy (welshfella):

@emily_98 - do you see how to get the values of b and c?

OpenStudy (emily_98):

message me

OpenStudy (welshfella):